# When Is Diversity Rewarded in Cooperative Multi-Agent Learning?

### v1

@@ I.3 $N=M=4$
where $\mathcal{N}=\{1,\ldots,n\}$ denotes the set of agents, $\mathcal{S}$ is the state space, and, $\left\{\mathcal{O}_{i}\right\}_{i\in\mathcal{N}}$ and $\left\{\mathcal{A}_{i}\right\}_{i\in\mathcal{N}}$ are the observation and action spaces, with $\mathcal{O}_{i}\subseteq\mathcal{S},\;\forall i\in\mathcal{N}$. Further, $\left\{\sigma_{i}^{\theta}\right\}_{i\in\mathcal{N}}$ and $\mathcal{R}^{\theta}$ are the agent observation and reward functions, such that $\sigma_{i}^{\theta}:\mathcal{S}\mapsto\mathcal{O}_{i}$, and, $\mathcal{R}^{\theta}:\mathcal{S}\times\left\{\mathcal{A}_{i}\right\}_{i\in\mathcal{N}}\mapsto\mathbb{R}$. $\mathcal{T}^{\theta}$ is the stochastic state transition model, defined as $\mathcal{T}^{\theta}:\mathcal{S}\times\left\{\mathcal{A}_{i}\right\}_{i\in\mathcal{N}}\mapsto\Delta\mathcal{S}$, which outputs the probability $\mathcal{T}^{\theta}(s^{t},\left\{a^{t}_{i}\right\}_{i\in\mathcal{N}},s^{t+1})$ of transitioning to state $s^{t+1}\in\mathcal{S}$ given the current state $s^{t}\in\mathcal{S}$ and actions $\left\{a^{t}_{i}\right\}_{i\in\mathcal{N}}$, with $a^{t}_{i}\in\mathcal{A}_{i}$. $\gamma$ is the discount factor. Finally, $s_{0}^{\theta}\in\mathcal{S}$ is a the initial environment state. A PDec-POMDP represents a set of traditional Dec-POMDPs [41], where the observation function, the transition function, the reward function, and the initial state are conditioned on parameters $\theta$. This formalism is similar to the concepts of Underspecified POMDP [36] and contextual MDP [49].
@@
Appendix L Limitations and Open Questions
@@
L.1 Theoretical scope
  • Inseparable tasks. This work considers task allocation problems where tasks are separable (i.e., the subtasks are independent), and that team reward can be factorized as an inner symmetric aggregator followed by an outer symmetric aggregator. Extending the theory to inseparable rewards, where performance on one task causally affects performance on another task (e.g., coverage with overlap penalty, or pairwise task constraints), is an open problem.
  • Beyond task-allocation RL domains. While the benchmark domains we study (matrix game, multi-goal capture) and the additional settings covered in App. D (Colonel Blotto, level-based foraging) all fit naturally into the task-allocation template, many notable RL domains—team autonomous driving, multi-robot manipulation—might not be representable within a task allocation framework. Our heterogeneity analysis does not directly apply to these settings, and extending our results to them is important for getting a complete picture of the benefits of heterogeneity.
L.2 Algorithmic assumptions
  • Differentiable simulation. HED requires $\nabla_{\theta}G^{\theta}(\bm{\pi})$, hence a simulator that is end-to-end differentiable and tractable to back-propagate through. Many realistic environments still rely on non-smooth physics or black-box generators, requiring us to modify HED for these settings. We note that there are good methods for learning environment parameters in non-differentiable settings, including the PAIRED algorithm [36] and the MARL-based environment design algorithm of [35], which can potentially be used to extend HED—but we leave such extensions to future work.
L.3 Open questions
@@
    Learning dynamics vs. reward structure. The theory predicts whether a given reward structure enables an advantage to heterogeneous teams, not whether a particular learning algorithm will learn in response to it. Our experiments suggest, empirically, that heterogeneous agents will, in practice, learn to exploit heterogeneous reward structures; but can a formal link be established between our reward structure insights and what reward the learning dynamics converge to in practice?

### v2

@@ 3 $r_{31}$ $r_{32}$ $r_{33}$
where $\mathcal{N}=\{1,\ldots,n\}$ denotes the set of agents, $\mathcal{S}$ is the state space, and, $\left\{\mathcal{O}_{i}\right\}_{i\in\mathcal{N}}$ and $\left\{\mathcal{A}_{i}\right\}_{i\in\mathcal{N}}$ are the observation and action spaces, with $\mathcal{O}_{i}\subseteq\mathcal{S},\;\forall i\in\mathcal{N}$. Further, $\left\{\sigma_{i}^{\theta}\right\}_{i\in\mathcal{N}}$ and $\mathcal{R}^{\theta}$ are the agent observation and reward functions, such that $\sigma_{i}^{\theta}:\mathcal{S}\mapsto\mathcal{O}_{i}$, and, $\mathcal{R}^{\theta}:\mathcal{S}\times\left\{\mathcal{A}_{i}\right\}_{i\in\mathcal{N}}\mapsto\mathbb{R}$. $\mathcal{T}^{\theta}$ is the stochastic state transition model, defined as $\mathcal{T}^{\theta}:\mathcal{S}\times\left\{\mathcal{A}_{i}\right\}_{i\in\mathcal{N}}\mapsto\Delta\mathcal{S}$, which outputs the probability $\mathcal{T}^{\theta}(s^{t},\left\{a^{t}_{i}\right\}_{i\in\mathcal{N}},s^{t+1})$ of transitioning to state $s^{t+1}\in\mathcal{S}$ given the current state $s^{t}\in\mathcal{S}$ and actions $\left\{a^{t}_{i}\right\}_{i\in\mathcal{N}}$, with $a^{t}_{i}\in\mathcal{A}_{i}$. $\gamma$ is the discount factor. Finally, $s_{0}^{\theta}\in\mathcal{S}$ is a the initial environment state. A PDec-POMDP represents a set of traditional Dec-POMDPs (Oliehoek et al., 2016), where the observation function, the transition function, the reward function, and the initial state are conditioned on parameters $\theta$. This formalism is similar to the concepts of Underspecified POMDP (Dennis et al., 2020) and contextual MDP (Modi et al., 2018).
@@
Appendix P Stability of Bilevel Optimization in HetGPS
The Heterogeneity Gain Parameter Search (HetGPS) algorithm employs a bilevel optimization framework to simultaneously optimize environment parameters and agent policies. This appendix discusses the structure of this optimization problem, its convergence properties, practical stability, and alternatives for non-differentiable environments.
P.1 HetGPS as a Stackelberg Game
HetGPS can be formalized as a Stackelberg game, a hierarchical optimization problem involving a leader and followers (Simaan & Cruz, 1973). In our setting:
  1. 1.
    The Leader is the environment designer (the outer loop of HetGPS), which aims to maximize the heterogeneity gain $HetGain^{\theta}$ by adjusting the environment parameters $\theta$.
  2. 2.
    The Followers are the homogeneous and heterogeneous multi-agent teams (the inner loop), which aim to maximize their respective returns $G^{\theta}(\pi)$ by optimizing their policies $\pi_{het}$ and $\pi_{hom}$ within the environment defined by $\theta$.
The leader’s objective function (the heterogeneity gain) depends on the optimized policies of the followers, which, in turn, depend on the parameters $\theta$ set by the leader. Formally, the objective is:
$\max_{\theta}\left[G^{\theta}(\pi^{*}_{\text{het}}(\theta))-G^{\theta}(\pi^{*}_{\text{hom}}(\theta))\right]$ (1)
where $\pi^{*}(\theta)$ represents the optimized policies for a given environment configuration $\theta$.
P.2 Convergence and Stability
Generally speaking, multi-agent reinforcement learning is a concurrent optimization process that faces non-stationarity as agents constantly adapt to one another’s evolving policies (Zhang et al., 2019). HetGPS extends this challenge as agents must also adapt to a changing environment. Consequently, formal convergence guarantees to a *global* optimum remain an open question with regards to HetGPS in particular, but also MARL algorithms in general. However, recent theoretical work in environment co-design has established conditions under which convergence of bilevel optimization processes similar to HetGPS to *local* optima can be guaranteed, such as requiring sufficient smoothness of the environment dynamics and policy updates (Gao et al., 2024).
Despite the theoretical complexities inherent in multi-agent learning and bilevel optimization, as shown in Sec. 5 and App. Q, HetGPS demonstrates strong empirical stability even under adversarial initializations. This stability is expected, as it mirrors the practical success observed in related co-design and automated curriculum learning literature (Dennis et al., 2020; Gao et al., 2024).
P.3 Advantage of Differentiable Simulation
In our experiments, HetGPS increased training time by roughly 25% compared to training agents in an environment with a fixed reward structure. Hence, it is highly efficient and does not impose much overhead. A key strength contributing to the efficiency of HetGPS is its use of differentiable simulation (e.g., VMAS (Bettini et al., 2022)). By leveraging backpropagation through the entire rollout, HetGPS computes the exact gradient $\nabla_{\theta}HetGain^{\theta}$. This approach is more sample-efficient than alternative methods that treat the environment design as a separate RL problem (e.g., PAIRED (Dennis et al., 2020) or Designer-RL (Gao et al., 2024; Amir et al., 2025)). Such methods rely on high-variance policy gradient estimates for the outer loop and often struggle with exploration inefficiency (Parker-Holder et al., 2021; Jiang et al., 2021; Xu et al., 2022). By utilizing exact gradients, HetGPS mitigates these issues.
P.4 Handling Non-Differentiable Environments
A requirement for the implementation of HetGPS presented in Alg. 1 is access to a differentiable simulator. When the environment involves non-smooth physics or black-box components, direct backpropagation is infeasible.
In such cases, the environment optimization step (Line 6 of Alg. 1) can be replaced with the gradient-free methods mentioned above, such as PAIRED (Dennis et al., 2020), the bilevel method from (Gao et al., 2024), or evolutionary strategies (Stanley et al., 2019). While these methods have empirically been shown to be stable and robust in other co-design settings, and may enable the extension of HetGPS to non-differentiable settings, they typically require more samples and may exhibit more noise compared to the direct backpropagation approach utilized in this work.
Appendix Q HetGPS Under Adversarial Initial Conditions
To evaluate the robustness of HetGPS to initialization, we repeated the Softmax experiment in Multi-Goal-Capture (Fig. 5a) with adverse initialization. We initialized the outer aggregator $U$ with $\tau=5$ (making it convex) and the inner aggregator $T$ with $\tau=-5$ (making it concave), which is the opposite of the concave-convex configuration predicted by theory to maximize heterogeneity gain.
As shown in Table 11, HetGPS successfully overcomes the adverse initialization and converges towards the theoretically optimal parameters (large positive $\tau$ for T, large negative $\tau$ for U).
Table 11: Convergence of HetGPS parameters ($\tau$) in the Softmax Multi-Goal-Capture experiment starting from adverse initialization ($\tau_{T}=-5,\tau_{U}=5$). Mean and standard deviation reported over 3 seeds.
Frames (M) 0 50 75 100
$\tau$ of T (Inner Agg.) $5.0\pm 0.0$ $13.32\pm 2.15$ $18.88\pm 3.96$ $22.95\pm 5.71$
$\tau$ of U (Outer Agg.) $5.0\pm 0.0$ $10.26\pm 1.70$ $14.59\pm 2.24$ $17.16\pm 2.43$
Appendix R Limitations and Open Questions
@@
R.1 Theoretical scope
  • Beyond task-allocation RL domains. The benchmark domains we study and the additional settings covered in App. E all fit into our abstract task-allocation framework: we can interpret agents’ state, such as goal proximity in Multi-goal-capture, or whether they captured an escaping agent in tag, abstractly as “efforts” $r_{ij}$ and represent the reward in terms of such efforts. This is what enables us to make predictions about these environments. Although our framework is quite general, and accommodates environments that one might not traditionally view as “task allocation” (such as football and tag), several notable multi-agent RL domains, e.g., multi-robot manipulation, might not be representable within this framework. Our heterogeneity analysis does not directly apply to these settings, and extending our results to them is important for getting a complete picture of the benefits of heterogeneity.
R.2 Algorithmic assumptions
  • Differentiable simulation. HetGPS requires $\nabla_{\theta}G^{\theta}(\bm{\pi})$, hence a simulator that is end-to-end differentiable and tractable to back-propagate through. We assume differentiability mainly for considerations of training efficiency. However, many realistic environments still rely on non-smooth physics or black-box generators, requiring us to modify HetGPS for these settings. We note that there are good, established methods for learning environment parameters in non-differentiable settings (at the cost of efficiency/increased noise). We discuss these in detail in App. P. However, we did not test such variants of HetGPS, and leave these extensions to future work.
R.3 Open questions
@@
    Learning dynamics vs. reward structure. The theory predicts whether a given reward structure enables an advantage to heterogeneous teams, not whether a particular learning algorithm will learn in response to it. This is connected to the difference between neural and behavioral heterogeneity that we emphasize throughout the paper. Our experiments suggest, empirically, that neurally heterogeneous agents will, in practice, learn to exploit heterogeneous reward structures (i.e., be behaviorally heterogeneous); but can a formal link be established between our reward structure insights and what reward the learning dynamics converge to in practice?